OpenStudy (anonymous):

Evaluate the expression for a = -1 and b = 5. show your work. 5a - 7b +b²

OpenStudy (fibonaccichick666):

what do you think you need to do to start?

OpenStudy (anonymous):

substitute a=-1 and b=5 into the equation

OpenStudy (anonymous):

5x-1 - 7x5 +5²??

OpenStudy (princeharryyy):

5*(-1) -7*5 + 5*5 = -5 - 35 +25 =-15

OpenStudy (fibonaccichick666):

ohoh ok they are multiplications ok yea so now, what do you do first by the order of operations?>

OpenStudy (anonymous):

no clue

OpenStudy (princeharryyy):

@AngelofVengeance6661

OpenStudy (fibonaccichick666):

ok so can you tell me the order of ops angel?

OpenStudy (anonymous):

no this is my brothers problem i sucked at algebra. and so does he

OpenStudy (fibonaccichick666):

ok so you need to check out the order of operations to begin this, do you remember PEMDAS or Please Excuse My Dear Aunt Sally?

OpenStudy (anonymous):

yeah a little

OpenStudy (fibonaccichick666):

if not it's ok, we can review

OpenStudy (anonymous):

parenthesis exponet multiply divide add subtract

OpenStudy (fibonaccichick666):

so give this a quick read for me, and one other favor, when you substitute put parentheses around the numbers you are inputting

OpenStudy (fibonaccichick666):

http://www.purplemath.com/modules/orderops.htm

OpenStudy (fibonaccichick666):

when you come back, tell me what you think you need to do first

OpenStudy (fibonaccichick666):

\[5(-1) - 7(5) +(5)^2\]

OpenStudy (anonymous):

im still lost

OpenStudy (anonymous):

should he show the parenthesis you just applied above?

OpenStudy (anonymous):

\[=-5-35+25\]

OpenStudy (anonymous):

=?

OpenStudy (anonymous):

55?

OpenStudy (anonymous):

solve it @AngelofVengeance6661

OpenStudy (anonymous):

nope

OpenStudy (fibonaccichick666):

yea, reason being is that it helps stop mistakes

OpenStudy (anonymous):

\[=-40+25\]

OpenStudy (fibonaccichick666):

so now, by order of operations, what will you do first?

OpenStudy (anonymous):

-15?

OpenStudy (fibonaccichick666):

just, work through with me and you'll getthe answer and understand it

OpenStudy (anonymous):

exponets?

OpenStudy (fibonaccichick666):

yes, so just baby steps here, can you simplify our exponent term \[5(-1) - 7(5) +(5)^2\]

OpenStudy (anonymous):

25

OpenStudy (fibonaccichick666):

so then, what are we left with?

OpenStudy (anonymous):

i think i may understand. im ocd so when it says parenthesis first and there arent any parenthesis i lose my mind completely

OpenStudy (fibonaccichick666):

ahh ok, well, there are invisible parentheses around the whole thing if it makes you feel better :) But anyways we have \[5(-1) - 7(5) +(5)^2=5(-1) - 7(5) +25\] right?

OpenStudy (anonymous):

5(−1)−7(5)+(5)2 -5 - 35 +25 -40 +25 = -15

OpenStudy (fibonaccichick666):

very good

OpenStudy (fibonaccichick666):

do the steps make sense?

OpenStudy (anonymous):

yes. thank you

OpenStudy (fibonaccichick666):

np, happy it makes sense. Thank you for wanting to learn. Please feel free to tag me when learining is the goal :)

OpenStudy (anonymous):

deffinitely will have a few more tonight.

OpenStudy (fibonaccichick666):

np if I'm awake, just tag me